OpenStudy (anonymous):

Write the expression (x+5)(x-2)(x+3) as a polynomial in standard form and state the number of solution. *PLEASE HELP ME STEP BY STEP*

OpenStudy (anonymous):

@pooja195

OpenStudy (campbell_st):

start with distributing (x + 5)(x - 2) the 2nd part to find the solutions... so all you need to do is solve the 2 linear factors x + 5 = 0, x-2 = 0 and x + 3 = 0

OpenStudy (anonymous):

-5,2,-3

OpenStudy (anonymous):

@campbell_st

pooja195 (pooja195):

Good so its got 3 solutions :)

OpenStudy (anonymous):

ok thank you